use long division to determine the quotient and remainder when f(x) = 4x³+6x²-2x+5 is divided by (2x-1)
1
Expert's answer
2020-11-26T15:24:42-0500

"\begin{matrix}
 & 2x^2 +4x+1 \ \ \ \ \ \ \ \ \ \\
 \ \ \ \ \ \ \ \ \ \ \ \ \ 2x-1 &)\overline{4x^3+6x^2-2x+5} \\
& - \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \\
& \underline{4x^3-2x^2} \ \ \ \ \ \ \ \ \ \ \ \ \ \\
& \ \ \ \ \ \ \ \ \ \ \ \ \ \ 8x^2-2x+5 \\
& \ \ - \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \\
& \ \ \ \ \ \ \ \ \underline{8x^2-4x} \\
& \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ 2x+5 \\
& \ \ - \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \\
& \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \underline{2x-1} \\
& \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ 6 \\
\end{matrix}"

Therefore,


"\dfrac{4x^3+6x^2-2x+5}{2x-1}=2x^2+4x+1+\dfrac{6}{2x-1}"

Quotient is "2x^2+4x+1."

Remainder is "6"